# LiDAR-BIND-T: Improved and Temporally Consistent Sensor Modality Translation and Fusion for Robotic Applications

### v2

@@ Abstract
This paper extends LiDAR-BIND, a modular multi-modal fusion framework that binds heterogeneous sensors (radar, sonar) to a LiDAR-defined latent space, with mechanisms that explicitly enforce temporal consistency. We introduce three contributions: (i) temporal embedding similarity that aligns consecutive latent representations, (ii) a motion-aligned transformation loss that matches displacement between predictions and ground truth LiDAR, and (iii) windowed temporal fusion using a specialised temporal module. We further update the model architecture to better preserve spatial structure. Evaluations on radar/sonar-to-LiDAR translation demonstrate improved temporal and spatial coherence, yielding lower absolute trajectory error and better occupancy map accuracy in Cartographer-based SLAM (Simultaneous Localisation and Mapping). We propose different metrics based on the Fréchet Video Motion Distance (FVMD) and a correlation-peak distance metric providing practical temporal quality indicators to evaluate SLAM performance. The proposed temporal LiDAR-BIND, or LiDAR-BIND-T, maintains plug-and-play modality fusion while substantially enhancing temporal stability, resulting in improved robustness and performance for downstream SLAM.
Index Terms:
Autonomous Vehicle Navigation, Deep Learning Methods, Robust/Adaptive Control, SLAM, Sensor Fusion
@@ I Introduction
Reliable perception of the surrounding environment is a fundamental requirement for autonomous systems, such as self-driving vehicles and mobile robots [1], [2]. Such systems typically employ a suite of optical sensors, namely cameras and LiDAR (Light Detection and Ranging), to construct a high-fidelity representation of their surroundings. However, while these sensors provide rich and dense data under ideal circumstances, their performance can be severely compromised by adverse optical environmental conditions like fog, rain, snow, or smoke. Consequently, a strong reliance on only these optical sensing modalities renders these systems vulnerable to critical perception errors and navigational failures, risking safety and operational reliability.
Given the shared vulnerabilities of optical sensors, we advocate the addition of other sensing modalities, such as mmWave radar and sonar, to increase robustness and accuracy in perception. Incorporating these diverse sensor types helps to reduce the range of conditions under which the system may fail, as these alternative modalities are often less susceptible to environmental factors that typically affect optical sensors. However, fusing data from such heterogeneous sensors remains challenging due to fundamental differences in characteristics such as spatial resolution, field of view, update rates and intrinsic modality properties. For instance, LiDAR provides precise and dense depth measurements in favourable optical conditions, but its performance deteriorates significantly in adverse optical environments. In contrast, radar is more robust under poor optical conditions, yet lacks the capability to deliver dense and high-resolution environmental measurements.
@@
Previous work proposed several techniques to increase the robustness of existing navigation frameworks by incorporating radar and sonar as additional modalities using our fusion methods, thereby significantly improving performance under adverse optical environmental conditions [4], [5]. By leveraging a deep learning-based approach, this method mapped measurements from one sensor modality to another, enabling the translation between modalities. In other words, the goal is to predict how one sensor modality should have perceived the environment based on the measurements acquired by a different sensor. Two distinct and separate modality translation models are proposed: one to predict 2D LiDAR data based on mmWave radar measurements and another to predict LiDAR data from ultrasound readings, employing the eRTIS sensor developed by our research group [6]. During deployment, the system can dynamically choose between using the original LiDAR data, which may be unreliable under adverse optical conditions, or the predictions generated by our models, which are comparatively more robust to degraded optical conditions. This decision-based fusion framework increases the overall robustness and safety of the perception system and enables effective operation in scenarios where LiDAR data alone would be insufficient. However, this selection-based approach limits information fusion by forcing a choice between sensor inputs at each timestep, preventing full use of complementary sensor strengths and potentially reducing performance in challenging conditions.
To address these limitations, the work in [3] proposed a learning-based framework that establishes a shared latent embedding space across different perception modalities, thereby enabling seamless sensor data fusion within this learned domain. The method employs dedicated encoders for each sensor type to transform modality-specific inputs into feature embeddings; a decoder is then trained to reconstruct 2D LiDAR data from this cross-modal embedding space. This framework, named LiDAR-BIND, highlighting its use of LiDAR data as a binding modality for establishing the shared embedding space, eliminates the need for traditional selection-based approaches.
@@
The challenge we address in this paper is to produce temporally consistent poses using SLAM and the (fused) LiDAR-BIND predictions. This problem can be linked to a well-recognised challenge in machine learning, particularly in the field of generative modelling, where predictions may often lack coherent evolution over time. Although state-of-the-art methods exist to address temporal inconsistency, applying them in this multimodal fusion context is not straightforward. These techniques are typically optimised or trained for real-world RGB images and, as a result, perform suboptimally when used with non-RGB sensor modalities.
In this paper, we propose improvements to the LiDAR-BIND framework that enable the fusion of multiple time steps across different sensor modalities, thereby improving both temporal and spatial consistency in predictions. These improvements contribute to the main goal of better performance in downstream tasks such as Simultaneous Localisation and Mapping (SLAM). The main contributions of this work are as follows:
@@
This paper is organised as follows. We start with a review of state-of-the-art approaches for temporal consistency in deep learning and their relevance to multi-modal sensor fusion in Section II. Next, Section III discusses the typical metrics for temporal consistency and their limitations for our non-RGB sensor measurement predictions. Section IV describes our hypothesis on the temporal consistency problem observed in LiDAR-BIND and the challenges of improving the temporal quality of predictions using the shared latent space. In Section V, we present the improvements to the LiDAR-BIND framework. Section VI details the experimental setup and results. Section VII provides an ablation study of the proposed components. Finally, Section VIII discusses the results and outlines directions for future work.
@@ II Consistency in Predictions: Inspiration from State-of-the-Art
Achieving temporal consistency in deep learning models is a central challenge across fields such as robotics, control systems and image and video generation. As many state-of-the-art methods have been developed in these fields, their core principles provide valuable inspiration for improving prediction consistency in our LiDAR-BIND multi-modal sensor fusion framework. However, direct application of these methods to the sparse and heterogeneous sensor data often encountered in robotics (e.g., radar, sonar, LiDAR), as presented in Figure 5, is typically non-trivial. Here, we highlight key state-of-the-art approaches and how their ideas inform our proposed improvements. We present these techniques in the categories we observed most prominently across these domains: (1) learning temporally consistent representations, and (2) efficient (long-term) temporal modelling.
@@
The recent advances in image and video generation further highlight the importance of temporally and spatially consistent representations. For example, MoVideo, presented in [12], improved the temporal quality of generated videos by proposing a motion-aware framework that explicitly takes into account video depth and optical flow. This approach, along with methods such as ConsistI2V [13] and LATTE (Latent Diffusion Transformer) [14], underscores the significance of both spatial and temporal modelling within the latent representation space, a technique we will incorporate in our proposed improvements. Additionally, the method presented in [15] introduces a motion consistency loss, further demonstrating the benefits of accounting for underlying motion in video generation.
In this work, we extend and adapt the LiDAR-BIND framework to achieve better spatio-temporal consistency within the embedding space. This is accomplished by encouraging similarity between embeddings of consecutive measurements and implementing architectural modifications, which are further detailed in Section V. Our approach is inspired by the objective of learning robust representations capable of capturing temporal dependencies. By aligning the embeddings of consecutive frames, we ensure that the model learns to produce embeddings that are similar, but not identical, for temporally adjacent frames, thereby achieving both temporal smoothness and sensitivity to actual changes in the environment.
@@ III Metrics for Temporal Consistency and Their Limitations
In our application, conventional evaluation metrics face significant limitations due to the sparse and heterogeneous characteristics of our data, as illustrated in Figure 5. Many of these standard tests rely on pre-trained image classification or object detection models for typical camera imagery. However, these approaches are not applicable to our context, where the data does not consist of standard images but rather originates from our other sensor modalities. Therefore, there is a clear need for domain-specific evaluation methods that can effectively assess the relevant aspects of consistency within our dataset. In this work, we present an attempt at creating metrics evaluating the temporal quality of our predictions for our specific use case of multi-modal sensor fusion for SLAM.
@@ IV LiDAR-BIND consistency problem
Beyond facilitating this multimodal fusion into a robust and unified embedding and measurement prediction, the mechanism enables on-the-fly anomaly detection by comparing the orientations of embeddings produced by each modality. If a particular sensor produces an embedding that notably deviates from those of other modalities, the fusion process can be dynamically adapted to rely more heavily on the reliable modalities, thereby further enhancing system robustness. Additional details regarding the LiDAR-BIND framework and the fusion mechanism can be found in our earlier work [24].
@@
We believe this issue can be addressed by incorporating methods that fuse information from multiple consecutive timesteps, while still keeping computational requirements low and retaining the benefits of our framework, namely efficient low-level fusion of sensor measurements. To this end, we draw inspiration from state-of-the-art techniques in deep learning designed to boost temporal embedding alignment. Additionally, we modify the architecture of the model to rely more on convolutions as linear learnable operators instead of the full rank matrix multiplications used in our previous work, which helps preserve spatial information and ensures that local features remain consistent throughout the encoding process.
Latent Space Design Considerations: Furthermore, it is important to note that the LiDAR-BIND shared latent space is entirely defined by the binding modality (LiDAR). We refer to this latent space as ”shared” because it provides a common ground for aligning embeddings from different modalities and time steps. However, this configuration means that the space is only optimised to capture features relevant to 2D LiDAR data. Although this should be sufficient for the decoder to function correctly, it can present challenges for the encoders of other modalities. These encoders may have difficulty extracting relevant features in relation to the desired 2D LiDAR prediction, particularly since the other modalities are sparser and more affected by sensor and environmental noise. Such noise inevitably degrades the quality of the fused representation.
@@
Examining these results, we observe that the radar-only latent space configuration outperforms the other configurations in terms of absolute positional error (APE). This suggests that, because the embedding space is specifically optimised for radar data, it generates more consistent predictions, which leads to superior SLAM performance, particularly in scenarios where scan matching is critical. Additionally, this experiment highlights a trade-off inherent to the LiDAR-BIND framework between its modularity and flexibility and the temporal consistency and SLAM accuracy of its predictions. With LiDAR-BIND, it is possible to add new modalities after deployment, simply by training an appropriate encoder, owing to its phased training approach. However, this flexibility may come at the expense of reduced temporal consistency and SLAM performance.
The latent space defined using just LiDAR data allows for the fusion of radar data, but does not perform as well as the radar-only latent space in our tests. The configuration in which both LiDAR and radar data are used to define the latent representation demonstrates that optimising the embedding space jointly for both modalities can improve performance. However, this improvement is accompanied by a loss of modularity and flexibility. Specifically, such a configuration would require all desired modalities to be present and synchronised in the training dataset. This constraint makes the emergent cross-modal alignment achieved by the LiDAR-BIND framework—for example, the alignment between sonar and radar—unfeasible in the combined latent space approach.
@@ V Temporal Improvements to LiDAR-BIND
Figure 5: Example of processed range-azimuth heatmaps from the LiDAR-BIND dataset. These heatmaps represent processed sensor data, where for each azimuth angle, the range information is encoded in a 2D grid format.
@@ V-A Embedding Similarity Between Consecutive Measurements
In the original LiDAR-BIND framework, we trained encoders to map different modalities into a shared latent space, ensuring, for example, that a sonar measurement of a scene would have a similar embedding to a LiDAR measurement of the same scene. Building on this concept, we extend the training objective to enforce high similarity between embeddings of consecutive measurements across time. This approach assumes that, under normal operating conditions, sensor measurements between time steps are similar, provided the time interval is not too large and that, for a vehicle moving at a certain velocity, the environment as perceived by the sensor will not change drastically between close time intervals. Figure 6 visualises this concept. As input to our models, we do not change from the original LiDAR-BIND framework and use processed range-azimuth heatmaps of the sensor data, an example of which can be seen in Figure 5. These heatmaps are generated from the raw sensor data and represent the range information in a 2D grid format, where each cell corresponds to a specific azimuth angle and range value.

### v3

@@ Abstract
This paper extends LiDAR-BIND, a modular multi-modal fusion framework that binds heterogeneous sensors (radar, sonar) to a LiDAR-defined latent space, with mechanisms that explicitly enforce temporal consistency. We introduce three contributions: (i) temporal embedding similarity that aligns consecutive latent representations, (ii) a motion-aligned transformation loss that matches displacement between predictions and ground truth LiDAR, and (iii) windowed temporal fusion using a specialised temporal module. We further update the model architecture to better preserve spatial structure. Evaluations on radar/sonar-to-LiDAR translation demonstrate improved temporal and spatial coherence, yielding lower absolute trajectory error and better occupancy map accuracy in Cartographer-based SLAM (Simultaneous Localisation and Mapping). We propose different metrics based on the Fréchet Video Motion Distance (FVMD) and a correlation-peak distance metric providing practical temporal quality indicators to evaluate SLAM performance. The proposed temporal LiDAR-BIND, or LiDAR-BIND-T, maintains modular modality fusion while substantially enhancing temporal stability, resulting in improved robustness and performance for downstream SLAM.
@@ I Introduction
Reliable perception of the surrounding environment is a fundamental requirement for autonomous systems, such as self-driving vehicles and mobile robots [1], [2]. These systems typically employ a suite of optical sensors, namely cameras and LiDAR (Light Detection and Ranging), to construct a high-fidelity representation of their surroundings. While these sensors provide rich and dense data under ideal circumstances, their performance can be severely compromised by adverse optical environmental conditions like fog, rain, snow, or smoke. Consequently, a strong reliance on only these optical sensing modalities renders these systems vulnerable to critical perception errors and navigational failures, risking safety and operational reliability.
Given the shared vulnerabilities of optical sensors, the integration of additional sensing modalities, such as mmWave radar and sonar, to increase the robustness and accuracy of perception systems is beneficial. Owing to the differences in their physical principles and wavelengths, the signals emitted by these sensors interact with the environment in different ways. By incorporating such diverse sensor types, the system’s operating range is broadened and the conditions under which failures may occur are significantly reduced, as these alternative modalities are generally less susceptible to environmental factors that typically affect optical sensors. However, fusing data from such heterogeneous sensors remains challenging due to fundamental differences in characteristics such as spatial resolution, field of view, update rates and intrinsic modality properties. For instance, LiDAR provides precise and dense depth measurements in favourable optical conditions, but its performance deteriorates significantly in adverse optical environments [3]. In contrast, radar is more robust under poor optical conditions, yet lacks the capability to deliver dense and high-resolution environmental measurements.
@@
Previous work proposed several techniques to increase the robustness of existing navigation frameworks by incorporating radar and sonar as additional modalities using novel fusion methods, thereby significantly improving performance under adverse optical environmental conditions [5], [6]. By leveraging a deep learning-based approach, this method mapped measurements from one sensor modality to another, enabling the translation between modalities. In other words, the goal is to predict how one sensor modality should have perceived the environment based on the measurements acquired by a different sensor. Two distinct and separate modality translation models were proposed: one to predict 2D LiDAR data based on mmWave radar measurements and another to predict LiDAR data from ultrasound readings, employing the eRTIS ultrasound sensor [7]. During deployment, the system can dynamically choose between using the original LiDAR data, which may be unreliable under adverse optical conditions, or the predictions generated by these models, which are comparatively more robust to degraded optical conditions. This selection-based fusion framework increases the overall robustness and safety of the perception system and enables effective operation in scenarios where LiDAR data alone would be insufficient. However, the selection-based approach limits information fusion by forcing a choice between sensor inputs at each timestep. Making the robustness improvements dependent on the algorithm or method choosing the input sensor at each timestep, prevents full use of complementary sensor strengths and potentially reduces performance in challenging conditions.
To address these limitations, the work in [4] proposed a learning-based framework that establishes a shared latent embedding space across different perception modalities, thereby enabling seamless sensor data fusion within this learned domain. The method employs dedicated encoders for each sensor type to transform modality-specific inputs into feature embeddings; a decoder is then trained to reconstruct 2D LiDAR data from this cross-modal embedding space. This framework, named LiDAR-BIND, highlighting its use of LiDAR data as a binding modality for establishing the shared embedding space, eliminates the need for selection-based approaches.
@@
The challenge we address in this paper is to produce temporally consistent poses using SLAM and the (fused) LiDAR-BIND predictions. This problem can be linked to a well-recognised challenge in machine learning, particularly in the field of generative modelling, where predictions may often lack coherent evolution over time. Although state-of-the-art methods exist to address temporal inconsistency, applying them in this multimodal fusion context is not straightforward, and has not been attempted before. These techniques are typically optimised or trained for real-world RGB images and, as a result, perform suboptimally when used with non-RGB sensor modalities.
In this paper, we propose improvements to the LiDAR-BIND framework that enable the fusion of multiple time steps across different sensor modalities, thereby improving both temporal and spatial consistency in predictions. These improvements to the predictions contribute to the main goal of better performance in downstream tasks such as Simultaneous Localisation and Mapping (SLAM). The main contributions of this work are as follows:
@@
This paper is organised as follows. We start with a review of state-of-the-art approaches for temporal consistency in deep learning and their relevance to multi-modal sensor fusion in Section II. Next, Section III discusses the typical metrics for temporal consistency and their limitations for the non-RGB sensor measurement predictions. Section IV describes our hypothesis on the temporal consistency problem observed in LiDAR-BIND and the challenges of improving the temporal quality of predictions using the shared latent space. In Section V, we present the improvements to the LiDAR-BIND framework. Section VI details the experimental setup and results. Section VII provides an ablation study of the proposed components. Finally, Section VIII discusses the results and outlines directions for future work.
@@ II Consistency in Predictions: Inspiration from State-of-the-Art
Achieving temporal consistency in deep learning models is a central challenge across fields such as robotics, control systems and image and video generation. As many state-of-the-art methods have been developed in these fields, their core principles provide valuable inspiration for improving prediction consistency in the LiDAR-BIND multi-modal sensor fusion framework. However, direct application of these methods to the sparse and heterogeneous sensor data often encountered in robotics (e.g., radar, sonar, LiDAR), as shown in Figure 9-a, is typically non-trivial. Here, we highlight key state-of-the-art approaches and how their ideas inform our proposed improvements. We present these techniques in the categories we observed most prominently across these domains: (1) learning temporally consistent representations, and (2) efficient (long-term) temporal modelling.
@@
The recent advances in image and video generation further highlight the importance of temporally and spatially consistent representations. For example, MoVideo, presented in [13], improved the temporal quality of generated videos by proposing a motion-aware framework that explicitly takes into account video depth and optical flow. This approach, along with methods such as ConsistI2V [14] and LATTE (Latent Diffusion Transformer) [15], underscores the significance of both spatial-temporal modelling within the latent representation space, a technique we will incorporate in our proposed improvements. Additionally, the method presented in [16] introduces a motion consistency loss, further demonstrating the benefits of accounting for underlying motion in video generation.
In this work, we extend and adapt the LiDAR-BIND framework to achieve better spatio-temporal consistency within the embedding space. This is accomplished by encouraging similarity between embeddings of consecutive measurements and implementing architectural modifications, which are further detailed in Section V. Our approach is inspired by the objective of learning robust representations capable of capturing temporal dependencies. By aligning the embeddings of consecutive frames, we ensure that the model learns to produce embeddings that are coherent, but not identical, for temporally adjacent frames, thereby achieving both temporal smoothness and sensitivity to actual changes in the environment.
@@ III Metrics for Temporal Consistency and Their Limitations
In our application, conventional evaluation metrics face significant limitations due to the sparse and heterogeneous characteristics of our data, as illustrated in Figure 9-a. Many of these standard tests rely on pre-trained image classification or object detection models for typical camera imagery. However, these approaches are not applicable to our context, where the data does not consist of standard images but rather originates from our other sensor modalities. Therefore, there is a clear need for domain-specific evaluation methods that can effectively assess the relevant aspects of consistency within our dataset. In this work, we present an attempt at creating metrics evaluating the temporal quality of our predictions for our specific use case of multi-modal sensor fusion for SLAM.
@@ IV LiDAR-BIND consistency problem
Beyond facilitating this multimodal fusion into a robust and unified embedding and measurement prediction, the mechanism enables on-the-fly anomaly detection by comparing the orientations of embeddings produced by each modality. If a particular sensor produces an embedding that notably deviates from those of other modalities, the fusion process can be dynamically adapted to rely more heavily on the reliable modalities, thereby further enhancing system robustness. Additional details regarding the LiDAR-BIND framework and the fusion mechanism can be found in earlier work [25].
@@
We believe this issue can be addressed by incorporating methods that fuse information from multiple consecutive timesteps, while still keeping computational requirements low and retaining the benefits of the framework, namely efficient low-level fusion of sensor measurements. To this end, we draw inspiration from state-of-the-art techniques in deep learning designed to boost temporal embedding alignment. Additionally, we modify the architecture of the model to rely more on convolutions as linear learnable operators instead of the full rank matrix multiplications used in the previous work, which helps preserve spatial information and ensures that local features remain consistent throughout the encoding process.
Latent Space Design Considerations: Next to this architectural issue, we also hypothesise that the modelling of the latent space itself contributes to the temporal inconsistency observed in LiDAR-BIND predictions. Currently this space is entirely defined by the binding modality (LiDAR), which does not suffer from the same temporal inconsistencies as the other modalities. For example multi-path reflections in both radar and sonar can cause significant variations in measurements between consecutive frames, even when the environment remains largely unchanged. This noise can lead to embeddings that are not only inconsistent over time but also misaligned with the corresponding LiDAR embeddings, which are more stable. We refer to this latent space as ”shared” because it provides a common ground for aligning embeddings from different modalities and time steps. However, this configuration means that the space is only optimised to capture features relevant to 2D LiDAR data. Although this should be sufficient for the decoder to function correctly, it can present challenges for the encoders of other modalities. These encoders may have difficulty extracting relevant features in relation to the desired 2D LiDAR prediction, particularly since the other modalities are sparser and more affected by sensor and environmental noise. Such noise inevitably degrades the quality of the fused representation.
@@
Examining these results, we observe that the radar-only latent space configuration outperforms the other configurations in terms of absolute positional error (APE). This suggests that, because the embedding space is specifically optimised for radar data, it generates more consistent predictions, which leads to superior SLAM performance, particularly in scenarios where scan matching is critical. We can interpret this as the model learning that different radar-scans, with and without noise (multi-path reflections), should still produce similar embeddings and predictions, as they represent the same scene. This results in more consistent predictions over time, which is crucial for effective SLAM performance. Additionally, this experiment highlights a trade-off inherent to the LiDAR-BIND framework between its modularity and flexibility and the temporal consistency and SLAM accuracy of its predictions. With LiDAR-BIND, it is possible to add new modalities after deployment, simply by training an appropriate encoder, owing to its phased training approach. However, this flexibility may come at the expense of reduced temporal consistency and SLAM performance.
The latent space defined using only LiDAR data allows the seamless fusion of radar and LiDAR in latent space; however, it does not perform as effectively as the radar-only latent space in our experiments. A likely explanation is that, where consecutive radar measurements often vary due to sensor noise, the corresponding LiDAR measurements remain relatively stable. Consequently, the model does not learn that such noise in the measurements should still produce similar embeddings and predictions. This leads to inconsistent predictions over time, thereby degrading SLAM performance. The configuration in which both LiDAR and radar data are used to define the latent representation demonstrates that optimising the embedding space jointly for both modalities can improve performance. However, this improvement is accompanied by a loss of modularity and flexibility. Specifically, such a configuration would require all desired modalities to be present and synchronised in the training dataset. This constraint makes the emergent cross-modal alignment achieved by the LiDAR-BIND framework—for example, the alignment between sonar and radar—unfeasible in the combined latent space approach.
@@
To summarize, we identified that the temporal consistency of LiDAR-BIND generated predictions can be significantly improved. The original framework, while effective at modality translation for individual timesteps, does not explicitly take into account the relationship between consecutive measurements. In addition, architectural and embedding space limitations were identified that lead to temporal discontinuities, particularly when translating from noisy sensor modalities like radar and sonar, where transient effects such as multi-path reflections can cause significant variations between frames. In the next section, we present our proposed improvements to the LiDAR-BIND framework that aim to address these challenges.
@@ V-A Embedding Similarity Between Consecutive Measurements
In the original LiDAR-BIND framework, encoders are trained to map different modalities into a shared latent space, ensuring, for example, that a sonar measurement of a scene would have a similar embedding to the synchronised LiDAR measurement of the same scene. Building on this concept, we extend the training objective to enforce high similarity between embeddings of consecutive measurements across time. This approach assumes that, under normal operating conditions, sensor measurements between time steps are similar, provided the time interval is not too large and that, for a vehicle moving at a certain velocity, the environment as perceived by the sensor will not change drastically between close time intervals. Figure 5 visualises this concept. As input to our models, we do not change from the original LiDAR-BIND framework and use processed range-azimuth heatmaps of the sensor data, an example of which can be seen in Figure 9-a. These heatmaps are generated from the raw sensor data and represent the range information in a 2D grid format, where each cell corresponds to a specific azimuth angle and range value.